40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Mon Oct 10 21:30:45 2011, Tara, Frank, Summary, Seismic, second leg replaced 
    Reply  Wed Oct 12 00:53:40 2011, Tara, Frank, Summary, Seismic, second leg replaced 
       Reply  Thu Oct 13 02:36:58 2011, Tara, Frank, Summary, Seismic, second leg replaced plot_2011_10_12.pngplot_2011_10_12.fig
          Reply  Fri Oct 14 14:40:51 2011, Tara, Frank, DailyProgress, Seismic, leakage in the legs, RFPD noise test beat_2011_10_13.figbeat_2011_10_13.pngerrnoise_volt.pngerrnoise.pngerrnoise.fig
Message ID: 705     Entry time: Thu Oct 13 02:36:58 2011     In reply to: 703     Reply to this: 706
Author: Tara, Frank 
Type: Summary 
Category: Seismic 
Subject: second leg replaced 

After optimized everything, I repeated the measurement that was done in this entry (noise at detection point). There is some improvement, the result is shown below.

 

==What have been done==

By "optimize everything", I meant:

  1. The Faraday Isolator was installed back in the setup, and optimized for maximum isolation. [add pic]
  2. Beam splitter and quarter wave plate sets (for double passed AOM and cavities) were optimized for minimum back reflection.
  3. EOM with half wave plate were adjusted for minimum RFAM (reduced by ~ 20 dB)
  4. Beam alignment to the cavities: visibility are up to 93% for both cavities now.

 Measurement Recap:  We want to check the noise from scattering noise or RFAM at the detection point, so we measure the noise at error point when the beam reflected off the cavity, or a mirror in front of the chamber (cf entry:700). We also want to see the dependent on power input, so we chose 10mW, 5mW and 1mW input power.

    Below are measurements from error point (Mixer out) which are calibrated to absolute frequency noise through the slope of error signal from each setup (power input of 10mW, 5mW and 1mW)

plot_2011_10_12.png

==Comments for the results==

  1.    The noise level when the beam reflected off the cavity goes down when we float the table and optimize everything ( red, green, blue are lower than pale pink (result from 2011_10_07).
  2.    The noise level when the beam reflected off the cavity do not change much with input power ( red, green, blue are about the same)
  3.    At 1mW the noise at high frequency (above 50Hz)  raises up for both cases (reflected off cavity / mirror) I'm not sure what happens. The calibration seems to be ok since the noise level at low frequency matches the results from 5mW and 10mW setup.
  4.    Seismic stack's resonant frequency at 6.7 Hz shows up more clearly after RFAM/ back reflection are minimized.

 

To Do: We will check the beat signal tomorrow. There should be improvement, since optics are optimized and the table is floated.

Attachment 2: plot_2011_10_12.fig  256 kB  Uploaded Thu Oct 13 03:39:49 2011
ELOG V3.1.3-